# MindFormer: A Transformer Architecture for Multi-Subject Brain Decoding via fMRI

### v1

@@ Abstract
Research efforts to understand neural signals have been ongoing for many years, with visual decoding from fMRI signals attracting considerable attention. Particularly, the advent of image diffusion models has advanced the reconstruction of images from fMRI data significantly. However, existing approaches often introduce inter- and intra- subject variations in the reconstructed images, which can compromise accuracy. To address current limitations in multi-subject brain decoding, we introduce a new Transformer architecture called MindFormer. This model is specifically designed to generate fMRI-conditioned feature vectors that can be used for conditioning Stable Diffusion model. More specifically, MindFormer incorporates two key innovations: 1) a novel training strategy based on the IP-Adapter to extract semantically meaningful features from fMRI signals, and 2) a subject specific token and linear layer that effectively capture individual differences in fMRI signals while synergistically combines multi subject fMRI data for training. Our experimental results demonstrate that Stable Diffusion, when integrated with MindFormer, produces semantically consistent images across different subjects. This capability significantly surpasses existing models in multi-subject brain decoding. Such advancements not only improve the accuracy of our reconstructions but also deepen our understanding of neural processing variations among individuals.
Figure 1: MindFormer architecture. The fMRI voxels obtained from observing the stimulus image are processed through the MindFormer to extract image features. These features are then utilized in conjunction with the Stable Diffusion model and a decoder to reconstruct the previously viewed image. The MindFormer is capable of distinguishing the origin of the information by using a learnable subject token, which allows it to identify from which specific subject the data was obtained.
@@ 1 Introduction
Brain decoding is the field focused on interpreting neural activity patterns to understand cognitive and sensory processes [1, 2, 3, 4, 5, 6]. Utilizing neuroimaging techniques such as functional magnetic resonance imaging (fMRI), researchers measure neural responses of brain activity to cognitive and sensory stimuli. Especially, visual brain decoding is a rapidly advancing field that aims to interpret brain activity to reconstruct visual experiences. Since the advent of deep learning, this field has significantly advanced [7, 8, 9, 10, 11, 12]. One popular approach to visual decoding involves mapping brain activity to latent spaces of generative models, such as generative adversarial networks (GANs) [13, 14, 15, 16]. With new large-scale fMRI datasets [17], recent advancements have also seen the use of diffusion models to improve reconstruction accuracy [18, 19, 20, 21, 22, 23, 24, 25]. The integration of diffusion models in brain decoding represents a significant leap forward, providing sophisticated tools to reconstruct and understand complex neural representations. However, challenges remain in achieving high fidelity, lightweight model and integrated-subject brain decoding.
In this work, we introduce a unified transformer architecture named MindFormer, which demonstrates exceptional performance in multi-subject brain decoding. MindFormer is specifically designed to generate semantically meaningful feature embeddings that facilitate fMRI-conditioned image generation using the Stable Diffusion model. We utilize the IP-adapter, as described in [26], to generate 16x768-dimensional feature embeddings from fMRI signals, which serve as conditions for the Stable Diffusion model, enhancing the decoding accuracy and reliability. Unlike previous studies that utilized CLIP embeddings, our semantic embedding via the IP-adapter is considerably smaller, reducing both computational costs and the risk of overfitting. Moreover, to synergistically combine training data from multiple subjects while accounting for individual differences, we introduce a subject-specific linear layer and the learnable subject token as inputs in the Transformer prompt. These subject specific components allow for the reconstruction of high-quality images from limited data sets by leveraging collective information from multiple subjects, improving its practical applicability in scenarios where data availability is restricted. Experimental results showed that our method maintains strong performance even with a limited amount of data by effectively leveraging the available common information across subjects to maximize accuracy and reliability.
@@
  • We developed a unified model called MindFormer, designed to generate semantically meaningful condition embedding modules for the Stable Diffusion model, specifically tailored for multi-subject brain decoding from fMRI signals.
  • MindFormer is a Transformer model that reduces dimensionality to decrease computational costs and minimize the risk of overfitting. Unlike previous methods that map brain signals into large CLIP image and text embeddings with dimensions of 257×768 and 77×768 respectively, MindFormer employs the IP-adapter to convert brain voxels into a compact 16x768-dimensional space. Additionally, it accounts for individual subject differences using a lightweight, subject-specific linear layer and learnable subject token, which significantly reduce the model’s size. These enhancements make MindFormer considerably more efficient than existing models.
  • To validate our method, we conducted experiments using the publicly available NSD dataset  [17]. We utilized a subset of 982 images that were commonly viewed by all four subjects as the test set. The training set consisted of 8,859 images, with low overlap between subjects. Experimental results confirm that the proposed method achieves excellent performance in multi-subject brain decoding.
@@ 2 Related Works
2.1 fMRI-to-Image Reconstruction Models
The fMRI-to-Image reconstruction models are the advanced approaches that aim to translate brain activity, captured through functional magnetic resonance imaging (fMRI), into visual images. These models learn complex mappings between neural signals and visual representations, enabling the generation of images that closely resemble the original stimuli perceived by the subjects. Since the advent of deep learning, the models leverages the deep learning frameworks, to interpret and reconstruct the visual experiences of subjects based on their neural activity patterns. The advancements have integrated generative models, such as Generative Adversarial Networks (GANs) [13, 14, 15, 16] and Variational Autoencoders (VAEs) [27], with fMRI data to improve the accuracy and quality of the reconstructed images. Seeliger et al. [16] explored the use of GANs for fMRI-to-image synthesis. Han et al.  [27] demonstrated the effectiveness of VAEs in reconstructing visual stimuli from brain activity. Advancements in deep learning have enabled the generation of not only natural scenes but also human faces  [28, 12] and video stimuli [29]. Moreover, advanced techniques like contrastive learning [30, 31] have been employed to align neural embeddings with visual embeddings more effectively. This alignment enhances the fidelity of the reconstructed images, ensuring that the generated visuals reflect the subjects’ visual experiences.
Figure 2: Training stage: The fMRI signals obtained after the $n$-th subject views an image are passed through a subject-specific linear layer. Subsequently, each signal is prepended with a learnable subject token and passed through the MindFormer Encoder. All subjects’ signals are processed through the same MindFormer Encoder. The resulting embeddings are then trained to match the image feature embeddings obtained from passing the images through the IP-Adapter. Inference Stage: These obtained embeddings are integrated into the stable diffusion process. The diffusion model utilizes these embeddings to iteratively denoise and reconstruct the image. The embeddings trained to match those generated by the IP-Adapter enhance the model’s ability to produce high-fidelity images that closely resemble the stimulus image.
@@ 2.2 Image Generation Diffusion Model
Another notable innovation in the field of brain decoding is the use of diffusion models. Diffusion models have gained popularity in generative modeling due to their capability to convert a noise vector into an output image via a reverse diffusion process [32, 33, 34]. Recent studies  [35, 36] have shown that diffusion models achieve superior image generation quality compared to GANs  [37, 38], further solidifying their importance in this field. These models have been particularly influential in advancing brain decoding, offering enhanced flexibility and precision in capturing the nuances of visual experiences encoded in brain activity  [18, 19, 20, 21, 22, 23, 24]. In most brain decoding studies that utilize diffusion models, researchers typically employ Stable Diffusion [39] or Versatile Diffusion [40] model. While Stable diffusion focuses on generating high-quality images by refining noise into coherent visuals, Versatile Diffusion facilitate substantial image variation, enabling tasks like style transfer. However, in Versatile Diffusion model, this variability can introduce challenges in brain decoding, as the reconstructed images may exhibit inconsistencies and artifacts. These discrepancies can complicate the accurate interpretation of neural activity, potentially reducing the fidelity of brain decoding outcomes.
Recent advancements in controllable image generation, such as ControlNet  [41] and T2I-adapter  [42], have shown that additional networks can guide image generation by plugging into existing text-to-image diffusion models. However, these methods often fall short in faithfully reproducing the reference image, primarily due to limitations in the cross-attention modules that inadequately merge image and text features. IP-Adapter  [26] addresses this issue by effectively integrating image features, enabling more accurate and detailed image generation based on reference inputs.
3 The MindFormer
@@ 3.1 Model Architecture
As shown in Fig. 1, the fMRI signals obtained after the $n$-th subject views an image are initially passed through a subject-specific linear layer. Following this, each signal is prepended with a unique learnable subject token and then processed through the MindFormer encoder. It is important to note that all subjects’ signals are processed through the same instance of the MindFormer encoder, ensuring a unified encoding process. The embeddings produced from this process are subsequently trained to match the image feature embeddings obtained when the images are passed through the IP-Adapter.
The central concept of the IP-Adapter revolves around its innovative decoupled cross-attention mechanism. Instead of employing a single cross-attention layer to handle both text and image features simultaneously, the IP-Adapter introduces a dedicated cross-attention layer specifically for image features. This separation enables the model to focus on learning more detailed and image-specific features, enhancing its ability to capture the unique characteristics of visual data. This training ensures that the neural representations align closely with the visual features embeddings of IP-Adapter, facilitating accurate and reliable image reconstruction from the fMRI signals. Consequently, the trained embedding, from the brain signal, are integrated into the Stable Diffusion process. The diffusion model utilizes these embeddings to iteratively denoise and reconstruct the image to generate semantically aligned images.
Specifically, as shown in Fig. 2, MindFormer comprises of the subject specific linear layer and a single transformer encoder that incorporates unique learnable subject token. The architecture of MindFormer encoder follows the Vision Transformer (ViT)  [43] encoder. Note that fMRI signals vary in size across subjects, primarily due to inherent differences in brain size and structure. To address the differing input voxel sizes, MindFormer maps each subject’s voxels ${\boldsymbol{v}}^{s}$ into a uniform dimension of 16×768 through individual linear layers $\mathcal{E}_{s}$ for each subject $s$. Then, in the position of BERT [44] and ViT [43]’s [Class] token, we prepend a learnable embedding token ${\boldsymbol{x}}_{subj}$ to the output ${\boldsymbol{x}}^{s}=\mathcal{E}_{s}({\boldsymbol{v}}^{s})$ of linear mapping. Addition use of learnable subject token ensures that the model accurately interprets and processes the neural data corresponding to multiple subject as well as each individual subject. Then, the position embeddings ${\boldsymbol{P}}$ are incorporated into the prepared embeddings to preserve positional information. The following steps proceed similarly to the transformer encoder in ViT. The Transformer encoder is composed of alternating layers of multi-headed self-attention (MSA) and MLP blocks. Layer normalization (LN) is applied before each block, with residual connections following each block. The MLP consists of two layers with a GELU activation function.
@@ 3.2 Training Objectives
Formally, we represent the 1D fMRI voxels from subject $s$ as ${\boldsymbol{v}}^{s}\in\mathbb{R}^{F_{s}}$, where $F_{s}$ denotes the subject specific size of the fMRI voxels. The corresponding image stimulus $I$ can be extracted as image feature embeddings ${\boldsymbol{E}}_{I}=[{\boldsymbol{e}}_{1},\cdots,{\boldsymbol{e}}_{N}]\in{\mathbb{R}}^{d\times N}$ using a pretrained IP-Adapter with $N=16$ and $d=768$. Additionally, MindFormer maps the brain voxels ${\boldsymbol{v}}^{s}$ into a 16×768-dimensional vector ${\boldsymbol{Z}}=[{\boldsymbol{z}}_{1},\cdots,{\boldsymbol{z}}_{N}]$, matching the dimension of the image feature embeddings ${\boldsymbol{E}}_{I}$ from the IP-Adapter  [26]. Then, MindFormer is trained with feature domain $l_{1}$-loss and contrastive learning loss between fMRI and Images:
@@
where $\alpha>0$ is a weight parameter. Specifically, the image feature-domain $l_{1}$ loss measures how MindFormer can predict the image feature ${\boldsymbol{E}}_{I}$ from IP-Adapter:
$\mathcal{L}_{1}({\boldsymbol{Z}},{\boldsymbol{E}}_{I})=\frac{1}{N}\sum_{i=1}^{N}\|{\boldsymbol{z}}_{i}-{\boldsymbol{e}}_{i}\|_{1}$ (2)
@@
$\mathcal{L}_{contrastive}({\boldsymbol{Z}},{\boldsymbol{E}}_{I})=\frac{1}{N}\sum_{i=1}^{N}\left(\textrm{log}\frac{e^{{\boldsymbol{z}}_{i}\cdot{\boldsymbol{e}}_{i}}}{\sum^{N}_{j=1}e^{{\boldsymbol{z}}_{i}\cdot{\boldsymbol{e}}_{j}}}\right)$ (3)
Figure 3: Natural scene and reconstructed image from human brain activity. All reconstructions are outputs from MindFormer. Additional reconstruction samples can be found in Figure 4 and Appendix A.2.
4 Experiments

### v2

# MindFormer: Semantic Alignment of Multi-Subject fMRI for Brain Decoding

@@ Abstract
Research efforts for visual decoding from fMRI signals have attracted considerable attention in research community. Still multi-subject fMRI decoding with one model has been considered intractable due to the drastic variations in fMRI signals between subjects and even within the same subject across different trials. To address current limitations in multi-subject brain decoding, here we introduce a novel semantic alignment method of multi-subject fMRI signals using so-called MindFormer. This model is specifically designed to generate fMRI-conditioned feature vectors that can be used for conditioning Stable Diffusion model for fMRI-to-image generation or large language model (LLM) for fMRI-to-text generation. More specifically, MindFormer incorporates two key innovations: 1) a subject specific token that effectively capture individual differences in fMRI signals while synergistically combines multi subject fMRI data for training, and 2) a novel feature embedding and training scheme based on the IP-Adapter to extract semantically meaningful features from fMRI signals. Our experimental results demonstrate that MindFormer generates semantically consistent images and text across different subjects. Since our MindFormer maintains semantic fidelity by fully utilizing the training data across different subjects by significantly surpassing existing models in multi-subject brain decoding, this may help deepening our understanding of neural processing variations among individuals.
Figure 1: Multi-subject brain decoding results by MindFormer. MindFormer can reconstruct semantically aligned images across subjects. Additional reconstruction samples can be found in Figure 4 and Appendix A.2.
Figure 2: MindFormer architecture. The fMRI voxels obtained from observing the stimulus image are processed through the MindFormer to extract image features. These features are then utilized in conjunction with the Stable Diffusion model and a decoder to reconstruct the previously viewed image. MindFormer is trained to counter the subject specific bias through learnable subject token.
@@ 1 Introduction
Brain decoding is a field dedicated to interpreting neural activity patterns to understand cognitive and sensory processes (Chen et al., 2014; Défossez et al., 2023; Du et al., 2022; Prince et al., 2022; Rao & Ballard, 1999; Schoenmakers et al., 2013). By utilizing neuroimaging techniques such as functional magnetic resonance imaging (fMRI), researchers measure brain activity in response to cognitive and sensory stimuli. A rapidly advancing area within this field is visual brain decoding, which aims to interpret neural signals to reconstruct visual experiences. The advent of deep learning has significantly propelled this field forward (Beliy et al., 2019; Gaziv et al., 2022; Gu et al., 2022; Horikawa & Kamitani, 2017; Shen et al., 2019; VanRullen & Reddy, 2019). One prevalent approach in visual decoding involves mapping neural activity to the latent spaces of generative models, such as generative adversarial networks (GANs) (Lin et al., 2022; Mozafari et al., 2020; Ozcelik et al., 2022; Seeliger et al., 2018). Recent advancements, fueled by new large-scale fMRI datasets (Allen et al., 2022), have seen the emergence of diffusion models, which enhance reconstruction accuracy (Chen et al., 2023; Lu et al., 2023; Mai & Zhang, 2023; Ozcelik & VanRullen, 2023; Scotti et al., 2024; Takagi & Nishimoto, 2023a; Xia et al., 2024; Wang et al., 2024). The integration of diffusion models in brain decoding marks a significant leap forward, providing advanced tools to reconstruct and interpret complex neural representations. Nonetheless, challenges remain in achieving high-fidelity reconstructions, lightweight models, and integrated subject-specific brain decoding.
In this work, we introduce a transformer-based multi-subject semantic alignment algorithm called MindFormer, which demonstrates exceptional performance in multi-subject brain decoding, particularly when combined with diffusion models or LLMs. MindFormer is specifically designed to generate semantically meaningful feature embeddings across multiple subjects to Stable Diffusion for image generation and LLMs for text generation. More specifically, to effectively integrate training data from multiple subjects while accounting for individual differences, we introduce a learnable subject token as inputs in the Transformer prompt. These components allow MindFormer to obtain semantically meaningful embedding even from limited datasets by leveraging collective information across subjects, improving its practical applicability in scenarios where data availability is restricted. Furthermore, we employ the IP-adapter, as described in Ye et al. (2023), to generate 16x768-dimensional feature embeddings from fMRI signals, which serve as conditioning inputs for the Stable Diffusion or LLMs. Unlike previous approaches that utilized CLIP embeddings, our use of the IP-adapter yields smaller, more efficient semantic embeddings, which reduces both computational costs and the risk of overfitting, and significantly enhances decoding accuracy and reliability. Experimental results demonstrate that our method maintains strong performance, even with limited data, by effectively utilizing shared information across subjects to maximize accuracy and reliability.
@@
  • We developed a semantic alignment method of multi-subject fMRI data using MindFormer to effectively integrate training data from multiple subjects while accounting for individual differences by using learnable subject token as inputs in the Transformer prompt. These components allow MindFormer to generate semantically meaningful condition embedding modules for the Stable Diffusion model, specifically tailored for multi-subject brain decoding from fMRI signals.
  • Unlike previous methods that map brain signals into large CLIP image and text embeddings with dimensions of 257×768 and 77×768, respectively, MindFormer utilizes the IP-adapter to transform brain voxels into a more compact 16×768-dimensional space. Along with a lightweight, subject-specific linear layer and a learnable subject token, the use of the IP-adapter significantly reduces the overall model size. These optimizations make MindFormer substantially more efficient than existing models.
  • To validate our method, we conducted experiments using the publicly available NSD dataset (Allen et al. (2022)). Experimental results confirm that the proposed method achieves excellent performance in multi-subject brain decoding. Also, our method effectively reconstructs high-quality images from limited datasets by leveraging shared information across subjects, maintaining strong performance even with constrained data availability.
  • We demonstrate the universality of MindFormer embedding by showing that its embedding can be used for LLM as inputs for accurate fMRI-to-text generation.
@@ 2 Related Works
2.1 $\mathrm{f}$MRI-to-Image Reconstruction Models
fMRI-to-image reconstruction models are advanced approaches designed to translate brain activity, captured through functional magnetic resonance imaging (fMRI), into visual images. These models learn complex mappings between neural signals and visual representations, allowing for the generation of images that closely resemble the original stimuli perceived by subjects. With the advent of deep learning, these models have increasingly leveraged deep learning frameworks to interpret and reconstruct visual experiences based on neural activity patterns. Recent advancements have integrated generative models, such as Generative Adversarial Networks (GANs) (Lin et al., 2022; Mozafari et al., 2020; Ozcelik et al., 2022; Seeliger et al., 2018) and Variational Autoencoders (VAEs) (Han et al., 2019), with fMRI data to improve the accuracy and quality of reconstructed images. For instance, Seeliger et al. (2018) explored the use of GANs for fMRI-to-image synthesis, while Han et al. (2019) demonstrated the effectiveness of VAEs in reconstructing visual stimuli from brain activity. Advances in deep learning have enabled the reconstruction of not only natural scenes but also human faces (Dado et al., 2022; VanRullen & Reddy, 2019) and video stimuli Wang et al. (2022). Additionally, techniques like contrastive learning (Chen et al., 2020; Radford et al., 2021) have been employed to better align neural embeddings with visual embeddings. This alignment significantly enhances the fidelity of the reconstructed images, ensuring that the generated visuals accurately reflect the subjects’ visual experiences.
Figure 3: Training stage: The fMRI signals from each subject are passed through a subject-specific linear layer. Subsequently, each signal is prepended with a learnable subject token and passed through the same MindFormer Encoder. The network is then trained to match the image feature embeddings obtained from passing the images through the IP-Adapter. Inference Stage: These obtained embeddings are integrated into the stable diffusion process as conditions. The diffusion model utilizes these embeddings to iteratively denoise and reconstruct the image.
@@ 2.2 Image Generation Diffusion Model
Another significant innovation in brain decoding is the use of diffusion models. Diffusion models have gained popularity in generative modeling due to their ability to transform noise vectors into output images through a reverse diffusion process (Ho et al., 2020; Nichol & Dhariwal, 2021; Song et al., 2020a). Recent studies (Dhariwal & Nichol, 2021; Song et al., 2020b) have demonstrated that diffusion models achieve superior image generation quality compared to GANs (Brock et al., 2018; Zhang et al., 2019), further establishing their importance in this field. These models have been particularly impactful in brain decoding, offering enhanced flexibility and precision in capturing the subtle nuances of visual experiences encoded in brain activity (Chen et al., 2023; Lu et al., 2023; Mai & Zhang, 2023; Ozcelik & VanRullen, 2023; Scotti et al., 2024; Takagi & Nishimoto, 2023a; Xia et al., 2024). In brain decoding studies that utilize diffusion models, researchers often employ the Stable Diffusion (Rombach et al., 2022) or Versatile Diffusion (Xu et al., 2023) models. Stable Diffusion focuses on generating high-quality images by refining noise into coherent visuals, while Versatile Diffusion enables substantial image variation, facilitating tasks like style transfer. However, this variability in Versatile Diffusion can introduce challenges in brain decoding, as the reconstructed images may display inconsistencies and artifacts. These discrepancies can complicate the accurate interpretation of neural activity, potentially reducing the fidelity of the decoding outcomes.
Recent advancements in controllable image generation, such as ControlNet (Zhang et al., 2023) and T2I-adapter (Mou et al., 2024), have shown that additional networks can guide image generation by plugging into existing text-to-image diffusion models. However, these methods often fall short in faithfully reproducing the reference image, primarily due to limitations in the cross-attention modules that inadequately merge image and text features. IP-Adapter (Ye et al., 2023) addresses this issue by effectively integrating image features, enabling more accurate and detailed image generation based on reference inputs. The central concept of the IP-Adapter revolves around its decoupled cross-attention mechanism. Instead of employing a single cross-attention layer to handle both text and image features simultaneously, the IP-Adapter introduces a dedicated cross-attention layer specifically for image features. This separation enables the model to focus on learning more detailed and image-specific features, enhancing its ability to capture the unique characteristics of visual data.
3 MindFormer
@@ 3.1 Model Architecture
As shown in Fig. 2, the fMRI signals obtained during the $n$-th subject’s viewing of an image are initially passed through a subject-specific linear layer. Following this, each signal is prepended with a unique learnable subject token and then processed through the MindFormer encoder. The embeddings produced from this process are subsequently trained to match the image feature embeddings obtained when the images are passed through the IP-Adapter. It is important to note that all subjects’ signals are processed through the same instance of the MindFormer encoder, ensuring a unified encoding process. For the case of image generation, the trained embedding, from the brain signal, are integrated into the Stable Diffusion process. The diffusion model utilizes these embeddings to iteratively denoise and reconstruct the image to generate semantically aligned images.
Specifically, as shown in Fig. 3, MindFormer comprises of the subject specific linear layer and a single transformer encoder that incorporates unique learnable subject token. The architecture of MindFormer encoder follows the Vision Transformer (ViT) (Dosovitskiy et al., 2020) encoder. Note that fMRI signals vary in size across subjects, primarily due to inherent differences in brain size and structure. To address the differing input voxel sizes, MindFormer maps each subject’s voxels ${\bm{v}}^{s}$ into a uniform dimension of 16×768 through individual linear layers $\mathcal{E}_{s}$ for each subject $s$. Then, in the position of BERT (Devlin et al., 2018) and ViT (Dosovitskiy et al., 2020)’s [Class] token, we prepend a learnable embedding token ${\bm{x}}_{subj}$ to the output ${\bm{x}}^{s}=\mathcal{E}_{s}({\bm{v}}^{s})$ of linear mapping. Addition use of learnable subject token is intended to decouple the individual bias of fMRI signal differences from the common representation across subjects, thereby allowing accurate interpretation of the neural data corresponding to multiple subject as well as each individual subject. Then, the position embeddings ${\bm{P}}$ are incorporated into the prepared embeddings to preserve positional information. The following steps proceed similarly to the transformer encoder in ViT. The Transformer encoder is composed of alternating layers of multi-headed self-attention (MSA) and MLP blocks. Layer normalization (LN) is applied before each block, with residual connections following each block. The MLP consists of two layers with a GELU activation function.
@@ 3.2 Training Objectives
Formally, we represent the 1D fMRI voxels from subject $s$ as ${\bm{v}}^{s}\in\mathbb{R}^{F_{s}}$, where $F_{s}$ denotes the subject specific size of the fMRI voxels. The corresponding image stimulus $I$ can be extracted as image feature embeddings $\mathbb{E}_{I}=[{\bm{e}}_{1},\cdots,{\bm{e}}_{N}]\in{\mathbb{R}}^{d\times N}$ using a pretrained IP-Adapter with $N=16$ and $d=768$. Additionally, MindFormer maps the brain voxels ${\bm{v}}^{s}$ into a 16×768-dimensional vector ${\bm{Z}}=[{\bm{z}}_{1},\cdots,{\bm{z}}_{N}]$, matching the dimension of the image feature embeddings $\mathbb{E}_{I}$ from the IP-Adapter (Ye et al., 2023). Then, MindFormer is trained with feature domain $l_{1}$-loss and contrastive learning loss between fMRI and Images:
@@
where $\alpha>0$ is a weight parameter. Specifically, the image feature-domain $l_{1}$ loss measures how MindFormer can predict the image feature $\mathbb{E}_{I}$ from IP-Adapter:
$\mathcal{L}_{1}({\bm{Z}},\mathbb{E}_{I})=\frac{1}{N}\sum_{i=1}^{N}\|{\bm{z}}_{i}-{\bm{e}}_{i}\|_{1}$ (2)
@@
$\mathcal{L}_{contrastive}({\bm{Z}},\mathbb{E}_{I})=\frac{1}{N}\sum_{i=1}^{N}\left(\textrm{log}\frac{e^{{\bm{z}}_{i}\cdot{\bm{e}}_{i}}}{\sum^{N}_{j=1}e^{{\bm{z}}_{i}\cdot{\bm{e}}_{j}}}\right)$ (3)
Method ##\## Models Low-level High-level PixCorr $\uparrow$ SSIM $\uparrow$ Alex(2) $\uparrow$ Alex(5) $\uparrow$ Incep $\uparrow$ CLIP $\uparrow$ EffNet-B $\downarrow$ SwAV $\downarrow$ Takagi et al. 4 -- -- 83.0% 83.0% 76.0% 77.0% -- -- Brain-Diffuser 4 .254 .356 94.2% 96.2% 87.2% 91.5% .775 .423 MindEye 4 .309 .323 94.7% 97.8% 93.8% 94.1% .645 .367 MindBridge (Single-) 4 .148 .259 86.9% 95.3% 92.2% 94.3% .713 .413 Ours (Single-) 4 .241 .352 93.5% 97.5% 93.5% 93.6% .659 .356 MindBridge 1 .151 .263 87.7% 95.5% 92.4% 94.7% .712 .418 Ours (Multi-) 1 .243 .345 93.5% 97.6% 94.4% 94.4% .648 .350
Table 1: Quantitative comparison of MindFormer’s decoding performance against other models. The metrics presented are averaged across the data from 4 subjects. Unlike other methods, which generally require a separate model for each subject, our approach and MindBridge consolidate the process into one model. Among them, our approach achieved superior results in all metrics except for the CLIP score. Bold: best, underline: second best.
4 Experiments Results
Experiment Settings. The proposed model is implemented in PyTorch. The single subject model is trained on one NVIDIA RTX 3090 GPU with a 24GB memory, and the multi subject model is trained on one NVIDIA RTX V100 with a 32GB memorys. Across all experiments, the batch size is set to 4 per GPU and the epoch size is 50. The learning rate was set as $3\times 10^{-4}$, and the moment parameters of the AdamW optimization algorithm (Loshchilov & Hutter, 2017) were set as $\beta_{1}=0.9,\beta_{2}=0.999$.
Dataset. To better understand the task at hand, we illustrate the data used in our study. For all experiments, we used the widely-adopted Natural Scenes Dataset (NSD) (Allen et al., 2022), a public fMRI dataset containing high-resolution 7-Tesla fMRI scans of brain responses from eight healthy adult subjects viewing natural scenes from the MS-COCO dataset (Lin et al., 2014). Following common practices (Mai & Zhang, 2023; Ozcelik & VanRullen, 2023; Scotti et al., 2024; Takagi & Nishimoto, 2023a; Wang et al., 2024), our research primarily uses data from four subjects (subj01, 02, 05, 07) who completed all scan sessions. Specifically, only a subset of data—982 images—was commonly viewed by all four subjects and used as the test set. The remaining data, comprising 8,859 distinct images viewed by each subject, were used as the training set, resulting in 24,980 training samples without averaging across repetitions, similar to the method used by previous research. We utilize the dataset, preprocessed by Scotti et al. (2024), which consists of flattened fMRI voxels within the brain volume space corresponding to the “nsdgeneral” brain region. This region, defined by the authors of  Allen et al. (2022), includes the subset of voxels that are most responsive to visual stimuli.